# PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue

### v1

@@ 3.5 Zero-shot prediction
3.6 Efficient Implementation
Efficient grouped query attention. We opt to use grouped query attention (GQA) [ainslie2023gqa] when doing cross attention with tile embeddings. This leads to significant memory savings during training by projecting tile embeddings to a much smaller embedding space as keys/values for cross attention. We use the official FlashAttention [dao2022flashattention, dao2023flashattention2] kernel from the flash-attn library which has native support for GQA rather than a naïve implementation that repeats key-value heads to match query heads which would negate the potential memory savings. The flash-attn kernel also supports sequence packing which allows us to avoid padding/unpadding before and after attention.
4 Experimental results and discussion
4.1 PRISM2 introduces slide-level diagnostic dialogue
Figure 4: This figure demonstrates the examples of PRISM2 dialogues. a A breast specimen with IDC. We show WSIs in the initial message for illustrative purposes, but for each example PRISM2 views all the slides of the specimen at once. b A bladder biopsy. Despite being trained on simple questions, the model is able to accurately respond to compound questions. c A uterus specimen. The model correctly identifies and characterizes the carcinoma and additional findings except it incorrectly predicts FIGO Grade 1 instead of Grade 2. d A liver specimen from a patient with metastatic colorectal cancer. The model is able to be given additional context like some patient history and specimen description and detects the carcinoma and correctly predicts its origin and grade, but incorrectly says there are no treatment related changes.
While PRISM has the ability to generate short clinical report summaries, PRISM2 adapts an LLM to interpret WSIs for diagnostic question-answering. This enables dialogue, given a set of WSIs as shown in Fig. 4. As the LLM (Phi-3 Mini [abdin2024phi]) is pretrained and instruction-tuned, we can see in Fig. 4b that it can generalize to compound questions, asking about multiple concepts in one sentence, even while the training data asks only about one concept per question. Asking the model targeted questions during training is intended to help it focus on all of the indications present in clinical reports rather than only on those with the most verbose descriptions. Furthermore, as we will see in the next section (Sec. 4.2), being able to ask specific questions enables simple and powerful zero-shot predictions.
As shown in Fig. 5 and Fig. 6, PRISM2 produces verbose descriptions of specimens. Furthermore, these descriptions remain consistent when looking only at the cropped regions that the model attends to the most. The model’s attention maps, taken from the slide encoder’s cross-attention layer, are shown for a single core of a WSI and each of its top regions of interest. We compute attention maps by summing up all the scores per tile across attention heads and queries and multiplying them by the $\ell_{2}$-norms of the corresponding values.
Figure 5: The slide encoder attention heatmap is shown for a core in a single slide of a lymphoma specimen, along with the diagnostic report generated by PRISM2. When cropping to each of the three top-attended regions, the PRISM2 report remains consistent with the evidence in the regions of interest.
Figure 6: The slide encoder attention heatmap is shown for a core in a single slide of a gastric specimen, along with the diagnostic report generated by PRISM2. When cropping to each of the three top-attended regions, the PRISM2 report remains consistent with the evidence in the regions of interest.
4.2 Dialogue simplifies zero-shot prediction
Model Method PanCancer PanCancer Rare Breast Subtyping GI Suite
TITAN Contrastive 0.728 (0.716, 0.738) 0.704 (0.687, 0.720) - -
PRISM Contrastive 0.793 (0.783, 0.803) 0.753 (0.737, 0.769) - -
PRISM2 Contrastive 0.719 (0.707, 0.730) 0.718 (0.701, 0.735) - -
Dialogue (YN) 0.874 (0.865, 0.882) 0.861 (0.848, 0.874) 0.874 (0.849, 0.897) 0.797 (0.781, 0.813)
Dialogue (DYN) 0.880 (0.871, 0.888) 0.872 (0.859, 0.885) 0.847 (0.822, 0.870) 0.815 (0.796, 0.834)
(a) Balanced accuracy for binary classification tasks using private MSKCC data.
Model Method BRCA NSCLC RCC
TITAN Contrastive 0.895 (0.870, 0.917) 0.930 (0.915, 0.946) 0.900 (0.882, 0.915)
PRISM Contrastive 0.883 (0.858, 0.906) 0.877 (0.858, 0.895) 0.631 (0.597, 0.666)
PRISM2 Contrastive 0.885 (0.863, 0.904) 0.888 (0.870, 0.906) 0.894 (0.867, 0.919)
Dialogue (OPN) 0.922 (0.901, 0.942) 0.925 (0.909, 0.941) 0.902 (0.879, 0.925)
(b) Balanced accuracy on multi-class classification tasks using public TCGA data.
Table 1: Zero-shot performance (balanced accuracy) for various classification tasks. 95% confidence intervals (CI) computed with 1000 iterations of bootstrapping are shown in parentheses. The best result per task (p $<$ 0.05) is in bold. (a) Dialogue-based yes/no (YN and DYN) question-answering substantially outperforms contrastive zero-shot prediction in the challenging PanCancer detection task, where it is challenging to assemble all relative cancer-positive and cancer-negative prompts for zero-shot evaluation. Defining the negative prompts is even more challenging for the binary classes in the Breast Subtyping and GI Suite datasets, so contrastive zero-shot evaluation for these is omitted. While both the YN and DYN use yes/no questions, DYN also includes a generated report as context. (b) Dialogue-based open-ended (OPN) question-answering is on par with TITAN contrastive zero-shot prediction on public TCGA multi-class classification.
Existing multi-modal pathology foundation models have demonstrated successful CLIP-style contrastive zero-shot classification [shaikovski2024prism, ding2024multimodal] on clinical tasks. Although including a contrastive training objective naturally endows vision-language models with the ability to perform zero-shot classification, such an approach also incurs several shortcomings. First, the optimality of a given text prompt is sensitive to the distribution of the training data. Second, diagnoses that cover a wide set of morphological features are difficult to represent by a small number of text prompts. For example, a benign diagnosis may be implied from a pathologist’s report by the absence of notable findings without being explicitly stated. Or, in the case of classifying invasive breast cancer, enumerating all the possible less severe diagnoses belonging to the noninvasive class (e.g. DCIS, LCIS, ADH, ALH) is cumbersome and inaccurate, as these indications can be present alongside invasive cancer. There may not exist a canonical text prompt for a given diagnosis.
Dialogue-based question-answering, using yes/no questions, is a better fit for identifying whether an indication is present or absent as no positive or negative sets of prompts need to be defined. Both contrastive and dialogue-based zero-shot predictions are evaluated in Tab. 1 and Supplementary Tab A.2.1. Indeed, whereas dialogue-based question answering (‘Dialogue (YN)’) achieves a balanced accuracy of 0.874 on all cancers and 0.861 on rare cancers, contrastive zero-shot with PRISM2 achieves only 0.719 and 0.718, respectively, comparable to TITAN’s 0.728 for all cancers and 0.704 for rare cancers. Including the diagnosis prediction of PRISM2 in the context (‘Dialogue (DYN)’) further boosts performance (0.874 to 0.880 balanced accuracy; 0.933 to 0.943 AUC) on all cancers and (0.861 to 0.872 balanced accuracy; 0.929 to 0.943 AUC on rare cancers). Contrastive zero-shot prediction works well on constrained multi-class subtyping tasks like TCGA BRCA, NSCLC, and RCC (Tab. 7(b) and Supplementary Tab 1(b)). Even so, dialogue-based zero-shot prediction performs as well or better.
The aforementioned challenges with contrastive zero-shot prediction are evident when analyzing pan-cancer contrastive zero shot performance. Using the prompts in Supplementary Tab. A.3.3 ’PanCancer’ to evaluate both PRISM and PRISM2, the former obtains a higher balanced accuracy (0.793 vs 0.719) and area under the curve (AUC) (0.896 vs. 0.805). However, when stratifying by tissue type, the inter- and intra-model variability in AUC demonstrates the brittle relationship between prompt selection and contrastive zero shot performance. In particular, we observe low performance on breast tissue for PRISM2 (0.622 AUC) even though the tissue represents 21% of the training samples. Thirty two percent of all breast false positives are caused by an erroneous match between the specimen’s image embedding and the “sarcoma” text prompt embedding—a word which is rarely used during training to describe breast cancer. This analysis motivates the hypothesis that contrastive zero shot performance depends precariously on the choice of prompt.
Indeed, we demonstrate that PRISM2 breast pan-cancer detection performance can be substantially improved just by changing the prompt selection to those in Supplementary Tab. A.3.3. We (I) align text prompts more closely with the training corpus language and (II) restrict our analysis to the breast cohort of the pan-cancer detection dataset. We craft a set of prompts which are stylistically in-distribution with respect to the PRISM2 training reports and delineate more closely between the positive diagnoses (IDC, ILC) and the negative ones (DCIS, LCIS, ADH, ALH, and benign). When evaluated against these new prompts, PRISM2 and PRISM breast AUC values are comparable (0.756 vs 0.798). This result supports the hypothesis that the CLIP-style zero shot capability depends on (and is hindered by) the ability to find the appropriate set of in-distribution text prompts.
4.3 Dialogue improves the feature representation for diagnostic tasks
In Tab. 2, we evaluate the quality of the base and diagnostic embeddings produced by PRISM2 by training linear classifiers on them across a variety of tasks, covering the detection and subtyping of cancers and precursors to cancers, as well as the detection of inflammatory and benign conditions and biomarkers (see A.1 for a description of evaluation tasks and data).
PRISM2 PRISM2 PRISM TITAN
Diagnostic Base
PanCancer 0.965 (0.961, 0.969) 0.952 (0.947, 0.958) 0.947 (0.942, 0.953) 0.931 (0.924, 0.937)
PanCancer Rare 0.953 (0.945, 0.960) 0.932 (0.922, 0.941) 0.931 (0.921, 0.940) 0.913 (0.902, 0.923)
Breast Subtyping 0.979 (0.976, 0.982) 0.976 (0.972, 0.979) 0.975 (0.971, 0.978) 0.923 (0.916, 0.929)
GI Suite 0.977 (0.974, 0.979) 0.978 (0.976, 0.980) 0.971 (0.968, 0.973) 0.965 (0.961, 0.968)
Biomarkers 0.845 (0.830, 0.859) 0.857 (0.844, 0.870) 0.843 (0.828, 0.857) 0.846 (0.829, 0.860)
TCGA BRCA 0.960 (0.938, 0.977) 0.958 (0.938, 0.976) 0.949 (0.927, 0.969) 0.956 (0.939, 0.970)
TCGA NSCLC 0.979 (0.970, 0.986) 0.980 (0.972, 0.987) 0.976 (0.968, 0.984) 0.981 (0.973, 0.989)
TCGA RCC 0.989 (0.984, 0.993) 0.991 (0.987, 0.994) 0.985 (0.979, 0.991) 0.992 (0.988, 0.995)
[…]
The diagnostic embedding significantly improves performance on the challenging PanCancer detection task, covering 16 different cancer tissue origins manifesting both as primary and metastatic cancer across diverse tissues. While the base embeddings of PRISM2 (0.952 AUC) outperform the PRISM embeddings (0.947 AUC), the difference is not statistically significant; whereas, the diagnostic embeddings (0.965 AUC) significantly (p $<$ 0.05) outperform all others. Both PRISM2 and PRISM embeddings substantially improve over TITAN (0.931 AUC). A similar trend is seen when focusing on the rare cancer tissues of origin, with PRISM2 diagnostic embeddings (0.953 AUC) significantly outperforming base embeddings (0.932 AUC), PRISM (0.947 AUC), and TITAN (0.913 AUC). This result demonstrates that the LLM can further improve diagnostic image understanding beyond the representation in the base embeddings learned via contrastive alignment. The diagnostic embeddings similarly produce the best result on breast subtyping (cancers and precursors to cancer) with an AUC of 0.979 as compared to 0.976 with the base embeddings, 0.975 with PRISM and 0.923 with TITAN; however, there is not enough statistical power to conclude a difference for this task with respect to the base embeddings. The low performance of TITAN on this task appears to be attributable to lower performance on precursors to cancer, as shown in Supplementary Tab. A.2.4.
Whereas diagnostic embeddings appear well-suited for cancer and precursor detection, they do not appear to be optimal for biomarker detection. Although there is low statistical power (due to the small sample sizes), we can see that the base embedding (overall 0.857 AUC) tends to outperform the diagnostic embedding (overall 0.845 AUC) in 9 out of 10 tasks and achieves the top score in 7 out of 10 tasks, as compared to one top score for PRISM and two top scores for TITAN (see Supplementary Tab. A.2.6). Recall that the diagnostic embeddings are extracted from the hidden state of the LLM which is trained to discuss these diagnostic findings but is not trained to discuss biomarkers. Indeed, the base embeddings trained with contrastive alignment may serve as lower level features that may be more readily transferable to out-of-distribution tasks.
PRISM2 significantly (p $<$ 0.05) outperforms both PRISM and TITAN on the gastrointestinal (GI) Suite, including the detection of cancers, precursors to cancer, inflammatory conditions, and benign indications (see Supplementary Tab. A.2.5 for more detailed results per class). As with breast subtyping the lower performance of TITAN (and in this case PRISM) is in large part related to greater difficulty detecting precursors to cancer or benign conditions.
On out of distribution TCGA tasks (BRCA, NSCLC, and RCC), PRISM2 performs on par with PRISM and TITAN, as shown in Tab. 2.
5 Discussion
PRISM2 advances the state-of-the-art in slide-level computational pathology foundation models by tackling several open problems. First, by scaling the dataset size to 2.3M WSIs and almost 700K clinical reports, PRISM2 is able to better generalize to diverse clinical tasks. Second, by introducing a large language model, PRISM2 introduces a conversational interface. Finally, by introducing the diagnostic and base embeddings, PRISM2 provides representations that are highly aligned with standard clinical tasks as described using natural language as well as a base embedding for use in emerging computational pathology applications such as biomarker quantification.
6 Acknowledgments
We gratefully thank Philip Rosenfield from Microsoft and Djamilia Dierov from Paige for their contributions in making this collaboration possible, Philippe Mathieu for distributed inference support, Mark Fleishman for data support, and Wayne Hendricks and Alexander van Eck for infrastructure support.
The results published here are in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga.
[…]
Appendix A Appendix
A.1 Evaluation tasks and data
To evaluate the quality of model embeddings, we train linear classifiers on a variety of tasks. Specifically, given a training, validation, and testing subset of data, we use the L-BFGS solver from scikit learn to train a linear classifier to convergence with an $\ell_{2}$ norm regularization term selected according to the lowest loss on the validation subset.
@@ A.1 Evaluation tasks and data
Pan-cancer Detection is an in-house evaluation dataset containing 22,932 H&E WSIs collected from 6,142 specimens. Samples are sourced in near-equal quantities from Memorial Sloan Kettering Cancer Center (49%) and global external sites (51%). Of the 16 cancers included in the dataset, seven are rare. The National Cancer Institute (NCI) defines rare cancers as those with an annual incidence in the United States below 15 people per 100,000. Complimenting the evaluation dataset, the training and validation sets together contain 89,417 slides across 40,402 specimens and follow a more natural distribution for the frequency of different cancers.
Breast Subtyping is an in-house dataset comprising 11,,774 breast H&E WSIs spanning 2407 specimens. Each specimen is associated with six non-mutually exclusive labels indicating the presence or absence of invasive lobular carcinoma (ILC), invasive ductal carcinoma (IDC), ductal carcinoma in situ (DCIS), lobular carcinoma in situ (LCIS), atypical ductal hyperpalasia (ADH), and atypical lobular hyperplasia (ALH). The dataset is divided at a specimen level into 8 folds with a 75/12.5/12.5 percentage split between train, tune, and test subsets.

### v2

@@ 4.5 Evaluation methods
Linear probe. Beyond the direct setting, model embeddings are also evaluated by training a linear classifier or linear regression. To do so, model embeddings are generated for training, validation, and testing sets of data. Linear classifiers are trained to convergence for each of 24 L2 regularization weights, with the best classifier selected on the validation set. For survival analysis, a Cox proportional hazards model is fitted to the training set for each of 24 elastic regularization weights (L1 and L2, with an L1:L2 ratio of 0.0001), with the best regressor selected on the validation set.
Report completion. We demonstrate the completion of pathology reports with the CAP worksheet for biopsies with invasive carcinoma of the breast [4]. The specimen procedure and laterality fields and the tumor position field are omitted from this analysis, as they constitute sample collection information rather than diagnostic findings derived from H&E stained slides. The mitotic rate measurement is also excluded as the model was not trained to count. Filling the fields in the CAP template is posed as multiple-choice QA for multiclass fields (eg. histologic type, histologic grade) or as yes-no QA for binary and multilabel fields (eg. presence of DCIS or its architectural pattern). The questions and detailed performance metrics are shown in Fig. 5. In total, the report can be represented with nine sets of fields, grouped under histologic type, histologic grade, DCIS, and “other” (lymphovascular invasion and microcalcifications), filled by asking 5 multiple-choice questions and 9 yes-no questions. We measure the adjusted mean recall (AMR). For multiclass, this is equivalent to adjusted balanced accuracy (adjusted such that random chance is scored as 0).
For direct prediction with PRISM2, we infer a prediction probability as the probability of the selected option or of ‘yes’ vs. ‘no’. This allows us to evaluate whether the model predictions can be used directly or if they require further calibration, as is often done with clinical-grade products. We evaluate both the AMR of direct prediction (observed) and the best possible AMR achieved by calibrating the probability (max). Calibration is per-class thresholding for binary (and multilabel) tasks, or per-class reweighting for multiclass tasks. We find the best thresholds by sweeping thresholds per class and the best reweighting with numerical optimization.
4.6 Comparison models
Several slide- or part-level embedding models have previously been reported and are used for comparison. We describe each briefly here. PRISM [29] is the previous model developed by the PRISM2 team. It differs from PRISM2 in that it is trained using few samples, with fewer types of text outputs and a different training algorithm. TITAN [13] is a slide-level embedding model that uses vision-language pretraining. TITAN aggregates CONCH [24] in a three-stage processing using both synthetic captions and medical reports and a training dataset of approximately 340 thousand whole slide images. COBRA [20] uses tile representations from four different foundation models to serve as feature space augmentations for a contrastive self-supervised learning approach. COBRA aggregates tiles using a state-space architecture and is trained using approximately 3 thousand slides. Prov-Gigapath [38] uses a vision-only approach to creating a slide-level model. Tile embeddings are aggregated using a masked auto-encoding approach and a training dataset of approximately 170 thousand whole slide images. Virchow2 Mean refers to a baseline where the Virchow2 [42] tile embeddings are naively aggregated by simply averaging them all together. Both PRISM2 and COBRA aggregate Virchow2 tile embeddings.
4.7 Evaluation datasets
@@
Pan-cancer Detection is an in-house evaluation dataset containing 22,932 H&E WSIs collected from 6,142 specimens. Samples are scanned at Memorial Sloan Kettering Cancer Center (MSK) and sourced in near-equal quantities from MSK (49%) and WSIs submitted to MSK for review from globally diverse external sites (51%). Of the 16 cancers included in the dataset, seven are rare. The National Cancer Institute (NCI) defines rare cancers as those with an annual incidence in the United States below 15 people per 100,000. Complimenting the evaluation dataset, the training and validation sets together contain 89,417 slides across 40,402 specimens and follow a more natural distribution for the frequency of different cancers. Note that the linear classifier trained on the pan-cancer training set is used to evaluate invasive cancer detection via linear probing on the prostate, breast, and breast lymph node product benchmarks below.
Prostate product benchmark. This dataset contains 2947 blocks (3327 slides) of prostate needle core biopsies. Labels for the blocks are extracted from synoptic reports collected at MSK. This dataset has been curated to evaluate the standalone performance of Paige Prostate Detect, which is a tissue-specific, clinical-grade model. We use this dataset to evaluate invasive cancer detection.
Breast product benchmark. This dataset contains 190 slides with invasive cancer and 1501 benign slides, labelled individually by a pathologist according to presence of atypical ductal hyperplasia (ADH), atypical lobular hyperplasia (ALH), lobular carcinoma in situ (LCIS), DCIS, IDC, ILC and/or other subtypes. This dataset has been curated to evaluate the standalone performance of Paige Breast, which is a tissue-specific, clinical-grade model. We group IDC and ILC into an invasive cancer label to evaluate invasive cancer detection.
Breast lymph node product benchmark. This dataset contains 458 lymph node slides with metastasized breast cancer, and 295 benign lymph node slides. Each slide has been labeled by a pathologist according to presence of invasive cancer, and largest tumor on the slide is measured to categorize the tumor into macrometastasis, micrometastasis, or infiltrating tumor cells. We group the categories together into an invasive cancer label to evaluate invasive cancer detection.
Breast Subtyping is an in-house dataset comprising 11,774 breast H&E WSIs spanning 2407 specimens. Each specimen is associated with six non-mutually exclusive labels indicating the presence or absence of invasive lobular carcinoma (ILC), invasive ductal carcinoma (IDC), ductal carcinoma in situ (DCIS), lobular carcinoma in situ (LCIS), atypical ductal hyperpalasia (ADH), and atypical lobular hyperplasia (ALH). The dataset is divided at a specimen level into 8 folds with a 75/12.5/12.5 percentage split between train, tune, and test subsets.